# Higher-Order Staircase Codes

### v1

@@ II-F Recovery of Classical Staircase Codes $(L=1,M=1)$
Consider taking $M=1$ with the optimal Golomb ruler of order $M+1=2$ and using the family of permutations given in Example 1 or in Example 2. We get that $d_{1}=1$ and that $\pi_{1}(i,j)=(j,i)$ or equivalently that $\Pi_{1}(B)=B^{\mathsf{T}}$ (the matrix transpose). Moreover, for $S>1$, the requirement $M\leq\mathsf{lpf}(S)$ holds trivially since $\mathsf{lpf}(S)\geq 2\geq M$. We therefore recover the classical staircase codes of [2] exactly. This motivates the following definition for a bona fide generalized staircase code:
@@
A generalized staircase code is a weak generalized staircase code as in Definition 1 with the added properties that the permutations form an $(M+1,S)$-net and that the ruler is an order $(M+1)$ Golomb ruler.
@@ III Extended-Hamming-Based Generalized Staircase Codes
Staircase codes and related product-like codes [3] typically use a $t$-error-correcting Bose–Chaudhuri–Hocquenghem (BCH) component code $\mathcal{C}$ to protect each coded bit with $M+1=2$ component codewords. This typically requires that $t\geq 3$ to achieve error floors below $10^{-15}$ [15] which imposes a limit on our ability to improve the energy-efficiency of such schemes since the power consumption of a BCH decoder circuit is roughly proportional to $t^{2}$ [16]. However, Proposition 1 suggests that for our generalized staircase codes which are scattering, we can reduce $t$ without increasing the error floor by compensating with a larger bit degree $M+1$. Accordingly, we consider the use of generalized staircase codes with single-error-correcting, i.e., $t=1$, components. Reasoning heuristically, we expect the resulting schemes to be more energy-efficient than those employing a smaller $M+1$ with a larger $t$ since a given correctable error pattern can be corrected with less energy assuming a $t^{2}$ power estimate for the correction of an equal share $t$ of the errors by each of the $M+1$ component codewords. We further anticipate improved performance–complexity–latency tradeoffs in a wider variety of contexts in light of recent works [17, 18] which accomplish such goals via new product-like codes with bit degree increased from two to three.
III-A LUT-Free Decoding of Systematic Hamming Codes
@@
$\begin{pmatrix}0&0&0&\dots&1\\ 0&0&0&\dots&1\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&\dots&1\\ 0&0&1&\dots&1\\ 0&1&0&\dots&1\\ 1&1&1&\dots&1\\ \end{pmatrix}\text{.}$ (13)
@@
Unfortunately, (13) must in general have its columns permuted in order to have a corresponding systematic generator matrix. This would then require that both decoding and generation of columns of the parity-check matrix is done by lookup table (LUT). Both of these operations must be done repeatedly during the iterative decoding of generalized staircase codes. It is thus useful for efficient software-based simulation, and potentially for hardware implementation, to be able to perform these operations as simply as in the case of (13) unpermuted. We accomplish this by means of algebraically-defined systematizing permutations.
@@
In particular, we interpret the column index set $[2^{r-1}]$ as the ring of integers modulo $2^{r-1}$, i.e., $\mathbb{Z}_{2^{r-1}}$. All odd integers are invertible modulo $2^{r-1}$ so that any odd $a\in\mathbb{Z}_{2^{r-1}}$ and any possibly non-odd $b\in\mathbb{Z}_{2^{r-1}}$ define an algebraic permutation $\tau\colon\mathbb{Z}_{2^{r-1}}\longrightarrow\mathbb{Z}_{2^{r-1}}$ where
$\tau(j)=a\cdot j+b\pmod{2^{r-1}}$
@@
$\tau^{-1}(j)=a^{-1}\cdot(j-b)\pmod{2^{r-1}}\text{.}$
In the context of software-based simulation, computing $\tau$ and $\tau^{-1}$ can be faster on a modern processor than performing the equivalent table lookup. By a computer search, values of $a$ and $b$ for each $(r-1)\in\{3,4,\dots,15,16\}$ were found such that the the parity-check matrix (13) permuted according to $\tau$ has a systematic generator matrix. The resulting values are provided in Table II. This means that both decoding as well as generating columns of the parity-check matrix can be done as simply as computing $\tau^{-1}$ and $\tau$. Adaptation to a code shortened in the first $s$ positions is accomplished by replacing $b$ with $b+a\cdot s$.
Finally, we report that, somewhat counter-intuitively, the choice of column permutation for the extended Hamming parity-check matrix has a slight but measurable impact on iterative decoding performance even in the absence of shortening. This is because the ordering of the columns impacts the spatial correlations between errors of odd weight greater than or equal to three and the consequent miscorrections. As a result, the use of the extended Hamming codes defined by Table II leads to slightly improved performance relative to systematizing permutations which preserve runs of naturally ordered columns. The same performance gain is also realized by a random choice of systematizing permutation which suggests that this gain is due to a decorrelation effect.
TABLE II: Systematizing algebraic permutations for extended Hamming codes
@@
III-B Simulations
We consider a component code $\mathcal{C}$ of length $(M+1)S$ and dimension $(M+1)S-r$ which is obtained by shortening a parent extended Hamming code of length $2^{r-1}$ in the first $2^{r-1}-(M+1)S$ positions where ${r-1=\lceil\log_{2}((M+1)S)\rceil}$. The parent codes are as defined by Table II. Furthermore, the permutations are taken to be those of Example 2 thus requiring that $M\leq\mathsf{lpf}(S)$, and the rulers are taken to be the optimal Golomb rulers from Table I. We then assume transmission across a binary symmetric channel (BSC) and perform sliding window decoding of $W$ consecutive blocks at a time where one iteration comprises decoding all constraints in the window consecutively similarly to [2]. To facilitate simulation of the waterfall regime where error propagation events dominate, we convert to a block code via a termination-like strategy in which a frame is formed from $F$ consecutive blocks. For the last $W$ of these $F$ blocks, the information portion is presumed to be zero and only the parity portion is transmitted with the encoder and decoder state being reset between consecutive frames. This results in a block code rate of
$R=\frac{(S-r)(F-W)}{S(F-W)+Wr}$
which approaches $R_{\mathsf{unterminated}}=1-r/S$ as $F$ tends to infinity.
TABLE III: Simulation parameters.
$S$ 669 409 307 179 47
$M$ 3 3 3 4 4
$r$ 13 12 12 11 9
$F$ 725 926 885 1634 912
$W$ (blocks) 21 21 21 36 48
$W$ (Mbits) 9.399 3.513 1.979 1.153 0.106
$R_{\mathsf{unterminated}}$ 0.98057 0.97066 0.96091 0.93855 0.80851
$R$ 0.98000 0.97000 0.96000 0.93725 0.80000
iterations 3 3 4 4 6
gap (dB) 0.585 0.650 0.750 0.950 1.850
input BER 9.86e-4 1.57e-3 2.09e-3 3.25e-3 1.05e-2
Figure 1: Simulation results for extended-Hamming-based generalized staircase codes with parameters given in Table III.
A highly-optimized software-based simulator for the BSC performance of the resulting generalized staircase codes was developed. The simulator is capable of achieving simulated throughputs of several gigabits per second per core of a modern multi-core consumer processor which allows for direct verification of sub-$10^{-15}$ error floors. This is enabled both by the simplicity of syndrome decoding of Hamming components and the fact that decoder behaviour depends only on the syndromes of the received component codewords. Performing Bernoulli sampling of errors for a BSC with crossover probability $\epsilon$ as geometrically-distributed jumps between error locations leads to roughly $\epsilon S^{2}(M+1)$ primitive operations being needed to update the decoder state in response to the reception of a new block. Estimating crudely and heuristically, one decoding iteration similarly takes roughly $\epsilon WS^{2}(M+1)$ primitive operations.
A few representative design examples are considered with parameters given in Table III and the corresponding simulation results given in Fig. 1 as bit error rate (BER) curves. We parameterize the BSC by the equivalent gap to the hard-decision Shannon limit under binary phase-shift keying (BPSK) transmission on an additive white Gaussian noise (AWGN) channel at the respective rates $R$ of the codes considered, rather than by the crossover probability. The last two rows of Table III provide the gap to the hard-decision Shannon limit along with the corresponding BSC crossover probability or input BER for the last point of each curve, i.e., the point of least BER. Where zero bit errors were measured, a dashed line is used in Fig. 1 to indicate the simulated operating point with the number of error-free bit transmissions also indicated. For all but the last code considered, we simulate the transmission of at least $10^{16}$ bits at the last plotted operating point, providing strong confidence that BERs are below $10^{-15}$. Relative to codes using two triple-error-correcting BCH components per bit in [3, 2, 15], the proposed codes perform roughly $0.2$ to $0.5$ dB worse in terms of the gap to the hard-decision Shannon limit as the code rate ranges from $0.98$ down to $0.8$. In return, we expect significantly reduced power consumption. We further see significant reductions in decoding memory thus latency.
IV Generalized Tiled Diagonal Zipper Codes with Higher Bit Degree ($L\geq 1$, $M\geq 1$)
@@
Tiled diagonal zipper codes [3] are a family of scattering generalizations of staircase codes for which the encoding memory approaches half of that of classical staircase codes while maintaining a variable node degree of $M+1=2$. In this section, we further generalize our scattering generalized staircase codes from Section II to both include this family and to extend it to arbitrary bit or variable degree $M+1$. This generalization has two uses. The first is that the encoding and decoding memory can be reduced while keeping all else essentially constant. The second is that it provides flexibility in hardware design by allowing the sidelength to be tuned at a constant code rate. We refer to the resulting scattering code family as higher-order staircase codes since it subsumes and extends the generalized staircase codes of Definition 6.

### v5

@@ II-F Recovery of Classical Staircase Codes $(L=1,M=1)$
Consider taking $M=1$ with the optimal Golomb ruler of order $M+1=2$ and using the family of permutations given in Example 1 or in Example 2. We get that $d_{1}=1$ and that $\pi_{1}(i,j)=(j,i)$ or, equivalently, that $\Pi_{1}(B)=B^{\mathsf{T}}$ (the matrix transpose). Moreover, for $S>1$, the requirement $M\leq\mathsf{lpf}(S)$ holds trivially since $\mathsf{lpf}(S)\geq 2\geq M$. We therefore recover the classical staircase codes of [2] exactly. This motivates the following definition for a bona fide generalized staircase code:
@@
A generalized staircase code is a weak generalized staircase code as in Definition 2 with the added properties that the permutations form an $(M+1,S)$-net and that the ruler is an order $M+1$ Golomb ruler.
@@ III Extended-Hamming-Based Generalized Staircase Codes
Staircase codes and related product-like codes [3] typically use a $t$-error-correcting Bose–Ray-Chaudhuri–Hocquenghem (BCH) component code $\mathcal{C}$ to protect each coded bit with $M+1=2$ component codewords. This typically requires that $t\geq 3$ to achieve error floors below $10^{-15}$ [19] which imposes a limit on our ability to improve the energy-efficiency of such schemes since the energy cost of a BCH decoding is roughly $\mathcal{O}(t^{2})$ [20]. However, Proposition 1 suggests that for our generalized staircase codes which are scattering, we can reduce $t$ without increasing the error floor by compensating with a larger bit degree $M+1$. Accordingly, we consider the use of generalized staircase codes with single-error-correcting, i.e., $t=1$, components. Reasoning heuristically, we expect the resulting schemes to be more energy-efficient than those employing a smaller $M+1$ with a larger $t$ since the energy cost per corrected bit would be $\mathcal{O}(t^{2}/t)=\mathcal{O}(t)$ with the lower bound of Proposition 1 held constant.
@@
$\begin{pmatrix}0&0&0&0&\cdots&1\\ 0&0&0&0&\cdots&1\\ \vdots&\vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\cdots&1\\ 0&0&1&1&\cdots&1\\ 0&1&0&1&\cdots&1\\ 1&1&1&1&\cdots&1\end{pmatrix}\text{.}$ (13)
@@
Unfortunately, (13) must in general have its columns permuted in order to have a corresponding systematic generator matrix. This would then require that both decoding and generation of columns of the parity-check matrix is done by lookup table (LUT). Both of these operations must be done repeatedly during the iterative decoding of generalized staircase codes. It is thus useful, both for efficient software-based simulation and for hardware implementation, to be able to perform these operations as simply as in the case of (13) unpermuted. We provide some solutions to this problem for software and hardware respectively in Sections III-B and III-C.
III-A Simulations
We consider a component code $\mathcal{C}$ of length $(M+1)S$ and dimension $(M+1)S-r$ which is obtained by shortening a parent extended Hamming code of length $2^{r-1}$ in the first $2^{r-1}-(M+1)S$ positions where ${r-1=\lceil\log_{2}((M+1)S)\rceil}$. The parent codes are as defined by Table V later. Furthermore, the permutations are taken to be those of Example 2 thus requiring that $M\leq\mathsf{lpf}(S)$, and the rulers are taken to be the optimal Golomb rulers from Table I. We then assume transmission across a binary symmetric channel (BSC) and perform sliding window decoding of $W$ consecutive blocks at a time where one iteration comprises decoding all constraints in the window consecutively and $I$ iterations are performed before advancing the window by one block similarly to [2]. To facilitate simulation of the waterfall regime where error propagation events dominate, we convert to a block code via a termination-like strategy in which a frame is formed from $F$ consecutive blocks. For the last $W$ of these $F$ blocks, the information portion is presumed to be zero and only the parity portion is transmitted with the encoder and decoder state being reset between consecutive frames. This results in a block code rate of
$R=\frac{(S-r)(F-W)}{S(F-W)+Wr}$
which approaches $R_{\mathsf{unterminated}}=1-r/S$ as $F\to\infty$. The reader is referred to [3] for alternative decoding and termination-like procedures for similar families of codes.
Figure 3: Simulation results for extended-Hamming-based generalized staircase codes with parameters given in Table II.
TABLE II: Simulation parameters for Fig. 3.
$S$ $M$ $F$ $W$ (Mbits) $R$ $I$ gap (dB) in. BER
669 3 725 21 (9.40) 0.980 3 0.585 9.86e-4
409 3 926 21 (3.51) 0.970 3 0.650 1.57e-3
307 3 885 21 (1.98) 0.960 4 0.750 2.09e-3
307 3 717 17 (1.60) 0.960 4 0.750 2.09e-3
179 4 1634 36 (1.15) 0.937 4 0.950 3.25e-3
179 4 1089 24 (0.77) 0.937 4 0.950 3.25e-3
47 4 912 48 (0.11) 0.800 6 1.850 1.05e-2
We developed a highly-optimized software simulator for the BSC performance of the resulting extended-Hamming-based generalized staircase codes. This simulator is available online [21] and is capable of achieving simulated throughputs of several gigabits per second per core of a modern multi-core consumer processor which allows for direct verification of sub-$10^{-15}$ error floors. This is enabled both by the simplicity of syndrome decoding of Hamming components and the fact that decoder behaviour depends only on the syndromes of the received component codewords. Therefore, simulation can be performed directly in syndrome domain with individual syndromes represented by individual integers. Performing Bernoulli sampling of errors for a BSC with crossover probability $\epsilon$ as geometrically-distributed jumps between error locations leads to roughly $\epsilon S^{2}(M+1)$ primitive operations being needed to update the decoder state in response to the reception of a new block. Estimating crudely and heuristically, one decoding iteration similarly takes roughly $\epsilon WS^{2}(M+1)$ primitive operations.
A few representative design examples are considered with parameters given in Table II and the corresponding simulation results given in Fig. 3 as bit error rate (BER) curves. We parameterize the BSC by the equivalent gap to the hard-decision Shannon limit under binary phase-shift keying (BPSK) transmission on an additive white Gaussian noise (AWGN) channel at the respective rates $R$ of the codes considered, rather than by the crossover probability. The last two rows of Table II provide the gap to the hard-decision Shannon limit along with the corresponding BSC crossover probability or input BER for the last point of each curve, i.e., the point of least BER. Where zero bit errors were measured, a dashed line is used in Fig. 3 to indicate the simulated operating point with the number of error-free bit transmissions also indicated. For all but the last code considered, we simulate the transmission of at least $10^{16}$ bits at the last plotted operating point, providing strong confidence that BERs are below $10^{-15}$.
Relative to codes using two triple-error-correcting BCH components per bit in [3, 2, 19], the example codes perform roughly $0.2$ to $0.5$ dB worse in terms of the gap to the hard-decision Shannon limit as the code rate ranges from $0.98$ down to $0.8$. In return, we get reduced complexity and latency. Some heuristic estimates of performance, latency, and complexity in comparison to classical staircase codes [2] are provided in Tables III and IV. In high-throughput contexts, latency is directly proportional to the number of bits in the decoding window $WS^{2}$. On the other hand, the number of component codewords in the decoding window is about $WS$. In practice, only a fraction of these must be decoded each iteration and we assume this fraction to be constant for simplicity. Moreover, we assume that for the relatively small values of $M$ considered, the cost of each component code decoding is dominated by the cost of invoking the component code decoder $t^{2}$ rather than the cost of updating the syndromes of codewords involved in each corrected bit. This leads to a heuristic complexity measure of $IWSt^{2}$.
TABLE III: Heuristic estimates of performance, latency, and complexity for generalized staircase codes with $R\approx 0.937$
staircase [2, 22] this work
gap to Shannon at 1e-15 0.56 dB 0.95 dB
latency: $WS^{2}$ 1.8e6 7.7e5
iterations: $I$ 3 4
decodings per iteration: $WS$ 3.6e3 4.3e3
cost per decoding: $t^{2}$ 9 1
complexity score: $IWSt^{2}$ 9.7e4 1.7e4
TABLE IV: Heuristic estimates of performance, latency, and complexity for generalized staircase codes with $R\approx 0.96$
staircase [2, 23] this work
gap to Shannon at 1e-15 0.49 dB 0.75 dB
latency: $WS^{2}$ 4e6 1.6e6
iterations: $I$ 4 4
decodings per iteration: $WS$ 5e3 5.2e3
cost per decoding: $t^{2}$ 9 1
complexity score: $IWSt^{2}$ 18e4 2e4
III-B LUT-Free Decoding of Systematic Hamming Codes with Affine Permutations in $\mathbb{Z}_{2^{r-1}}$
We now revisit the problem of constructing systematic Hamming codes that admit a LUT-free decoding. This amounts to finding a permutation of the columns of (13) such that a corresponding systematic generator matrix exists and both the forward and inverse permutations are easily calculated without a LUT. Note that the existence of a systematic generator matrix is equivalent to the invertibility of the $r\times r$ matrix formed by the last $r$ columns of the parity-check matrix. We will refer to a permutation of the columns of (13) achieving this as a systematizing permutation.
@@ III-B LUT-Free Decoding of Systematic Hamming Codes with Affine Permutations in $\mathbb{Z}_{2^{r-1}}$
A natural approach to this problem is to consider affine permutations in $\mathbb{Z}_{2^{r-1}}$. In particular, we interpret the column index set $[2^{r-1}]$ as the ring of integers modulo $2^{r-1}$, i.e., $\mathbb{Z}_{2^{r-1}}$. All odd integers are invertible modulo $2^{r-1}$ so that any odd $a\in\mathbb{Z}_{2^{r-1}}$ and any possibly non-odd $b\in\mathbb{Z}_{2^{r-1}}$ define an algebraic permutation $\tau\colon\mathbb{Z}_{2^{r-1}}\longrightarrow\mathbb{Z}_{2^{r-1}}$ where
$\tau(j)=aj+b\pmod{2^{r-1}}$
@@
$\tau^{-1}(j)=a^{-1}(j-b)\pmod{2^{r-1}}\text{.}$
In the context of software-based simulation, computing $\tau$ and $\tau^{-1}$ can be faster on a modern processor than performing the equivalent table lookup. By a computer search, values of $a$ and $b$ for each $(r-1)\in\{3,4,\dots,16\}$ were found such that $\tau$ is systematizing. The resulting values are provided in Table V. This means that both decoding a syndrome as well as generating columns of the parity-check matrix can be done as simply as computing $\tau^{-1}$ and $\tau$. Adaptation to a code shortened in the first $s$ positions is accomplished by replacing $b$ with $b+as$.
We report that, somewhat counter-intuitively, the choice of column permutation for the extended Hamming parity-check matrix leads to a slight but measurable improvement in iterative decoding performance even in the absence of shortening. This is because the ordering of the columns impacts the spatial correlations between errors of odd weight greater than or equal to three and the consequent miscorrections. As a result, the use of the extended Hamming codes defined by Table V leads to slightly improved performance relative to systematizing permutations which preserve runs of naturally ordered columns. The same performance gain is also realized by a random choice of systematizing permutation which suggests that this gain is due to a decorrelation effect.
TABLE V: Parameters for systematizing affine permutations for extended Hamming codes
@@
III-C LUT-Free Decoding of Systematic Hamming Codes with Boolean Functions
TABLE VI: Parameters for systematizing permutations for extended Hamming Codes via Boolean Functions
$i$ $d^{(i)}$ $\sum_{u=0}^{d^{(i)}-1}s^{(i)}_{u}2^{u}$
2 2 0
3 2 0
4 3 0
5 3 1
6 3 2
7 2 0
8 4 0
9 4 1
10 4 2
11 4 3
12 4 4
13 4 5
14 4 6
15 3 0
In the context of software implementation, the approach presented in Section III-B is as good as we can hope for since modern processors already have dedicated multipliers. On the other hand, in the context of hardware implementation, we might wish to avoid realizing multipliers—which are expensive—altogether and directly seek a permutation calculable with a small number of logic gates. To this end, we extend an approach to LUT-free Hamming decoding provided in [24].
Firstly, we interpret the column index set $[2^{r-1}]$ as the vector space $\mathbb{F}_{2}^{r-1}$ taking $j\in[2^{r-1}]$ with binary representation $j=\sum_{u=0}^{r-2}b_{u}2^{u}$ with $b_{u}\in\{0,1\}$ to correspond to the binary vector $(b_{0},b_{1},\ldots,b_{r-2})\in\mathbb{F}_{2}^{r-1}$.
Consider the class of permutations of $\mathbb{F}_{2}^{r-1}$ which map $(b_{0},b_{1},\dots,b_{r-2})\in\mathbb{F}_{2}^{r-1}$ to $(c_{0},c_{1},\dots,c_{r-2})\in\mathbb{F}_{2}^{r-1}$ and take the form
$b_{0}$ $\mapsto c_{0}=b_{0}$
$b_{1}$ $\mapsto c_{1}=b_{1}$
$b_{i}$ $\mapsto c_{i}=b_{i}+\prod_{u=0}^{d^{(i)}-1}(s^{(i)}_{u}+b^{(i)}_{u})\;\text{for $2\leq i\leq r-2$}$
where $d^{(i)}$ and $(s^{(i)}_{0},s^{(i)}_{1},\dots,s^{(i)}_{d^{(i)}-1})$ for $22\leq i\leq r-$ 2 are parameters defining the permutation. We further impose that $d^{(i)}\leq i$ which guarantees invertibility since the expression
$c_{i}=b_{i}+\prod_{u=0}^{d^{(i)}-1}(s^{(i)}_{u}+b^{(i)}_{u})$
allows the recovery of $b_{i}$ from $c_{i}$ and $b_{0},b_{1},\dots,b_{i-1}$ and thus all $b_{i}$ from $c_{i}$ by induction. We will represent $(s^{(i)}_{0},s^{(i)}_{1},\dots,s^{(i)}_{d^{(i)}-1})$ with the corresponding integer $\sum_{u=0}^{d^{(i)}-1}s^{(i)}_{u}2^{u}$ for convenience.
This class of permutations generalizes the specific example given in [24] for $r=8$ and is easy to implement in hardware requiring for each $i$ a $d^{(i)}$-input AND gate, one XOR gate, and a NOT gate for every nonzero value of $s^{(i)}_{u}$.
Table VI defines a special choice of systematizing permutations for extended Hamming codes of length up to $2^{16}$ with the property that the table can be truncated at any row while maintaining the systematizing property. These were found by a computer search. For example, for extended Hamming codes of length $642^{6}=$ 64, the map defined by Table VI is
$(b_{0},b_{1},b_{2},b_{3},b_{4},b_{5})\mapsto\\ (b_{0},b_{1},b_{2}+b_{0}b_{1},b_{3}+b_{0}b_{1},b_{4}+b_{0}b_{1}b_{2},b_{5}+(1+b_{0})b_{1}b_{2})$
The last seven columns of our parity-check matrix are are the image of $\{57,58,\dots,63\}$ under this map evaluated over $\mathbb{F}_{2}^{6}$ with an all-ones row appended. This yields
$\begin{pmatrix}1&1&1&1&1&0&1\\ 1&1&1&1&1&1&0\\ 1&1&0&1&1&1&0\\ 0&0&1&1&1&1&0\\ 0&1&1&0&0&1&1\\ 1&0&1&0&1&0&1\\ 1&1&1&1&1&1&1\end{pmatrix}$
which is invertible. Example code is also provided online [25].
IV Higher-Order Staircase Codes ($L\geq 1$, $M\geq 1$)
@@
Tiled diagonal zipper codes [11] are a family of scattering generalizations of staircase codes for which the encoding and decoding memory approaches half of that of classical staircase codes while maintaining a variable node degree of $M+1=2$. In this section, we further generalize our scattering generalized staircase codes from Section II to both include this family and to extend it to arbitrary bit or variable node degree $M+1$. This generalization has two uses: The first is that the encoding and decoding memory can be reduced while keeping all else essentially constant. The second is that it provides flexibility in hardware designs by allowing the sidelength to be tuned while keeping all else essentially constant. The sidelength is a proxy for the degree of parallelism that is naturally admitted by an encoding or decoding circuit. We refer to the resulting scattering code family as higher-order staircase codes since it subsumes and extends the generalized staircase codes of Definition 7.